# Evaluating Named Entity Recognition: Comparative Analysis of Mono- and Multilingual Transformer Models on Brazilian Corporate Earnings Call Transcriptions

### v1

@@ 2.2 T5-based models: PTT5 and mT5
The Text-to-Text Transfer Transformer (T5) model is based on the Transformer encoder-decoder architecture and uses SentencePiece as its tokenizer Raffel2020 . To use T5, we need to map our problem to a text-to-text format, i.e. taking text as input and generating text as output Raffel2020 . T5 was pre-trained using “Colossal Clean Crawled Corpus” (C4) and it was evaluated on a wide variety of NLP tasks such as question answering, document summarization, and sentiment classification Raffel2020 . Instead of using a fixed embedding for each position like BERT, T5 uses relative position embeddings.
@@
Similar to BERT, T5 has also served as a base for other models. Carmo et al. Carmo2020 introduced, in 2020, a model based on T5 architecture and pre-trained from scratch using BrWaC (Brazilian Web as Corpus) dataset. The authors named this model PTT5 and released three versions on Hugging Face: Small (60M parameters), Base (220M parameters), and Large (740M parameters). They also recommend the use of the Base version.
@@
A multilingual version of T5 was introduced by Xue et al. Xue2021 in 2021. The model, named mT5, was pre-trained using a multilingual version of C4 (mC4 - 101 languages). mT5 was released in five versions: mT5-Small (300M parameters), mT5-Base (580M parameters), mT5-Large (1.2B parameters), mT5-XL (3.7B parameters), and mT5-XXL (13B parameters).
@@
T5 models can process longer inputs than BERT models. However, longer inputs may still lead to increased computational cost and memory usage during training and inference. We confirmed that PTT5 and mT5 provided on Hugging Face have an input of 512 length.
3 Related Works
Named Entity Recognition (NER) has been studied over the years using various datasets, label sets, and strategies. In 2015, Alvarado, Verspoor, and Baldwin Alvarado2015 constructed a NER dataset utilizing the CoNLL-2003 dataset available in the literature, in addition to a supplementary dataset created by them, which comprised domain-specific English texts. This supplementary dataset included texts extracted from financial agreements, sourced from 8 documents (totaling 54,256 words) randomly selected for manual annotation. They annotated four types of entities provided in the CoNLL-2003 dataset: LOCATION (LOC), ORGANIZATION (ORG), PERSON (PER), and MISCELLANEOUS (MISC). Furthermore, they automatically annotated all instances of “lender” and “borrower” as PERSON (PER). Experiments were conducted using various combinations of training and testing datasets, employing Conditional Random Fields (CRFs) for annotation. The highest F-score (82.7%) was attained when the CRF was trained and tested with the domain-specific dataset.
In 2019, Francis, Van Landeghem, and Moens Francis2019 developed a dataset using 3000 invoices from service providers such as insurance, telecommunications, banking, and tax companies. They employed optical character recognition (OCR) to extract English texts related to International Bank Account Number (IBAN) of the beneficiary, invoice number, invoice and due dates, total inclusive amount, total exclusive amount, total Value Added Tax (VAT) amount, structured communication, company name, company address, and other related fields required for invoice processing. The NER task was performed using an architecture consisting of an input layer, two stacked Bidirectional Long Short-Term Memory (Bi-LSTM), and a CRF layer. Due to variances in entity structure between the gold dataset [^2] A gold dataset is considered to be of high quality and accuracy due to the meticulous manual annotation process conducted by human annotators. and the extracted entities, the authors evaluated their strategy manually.
Two years later, in 2021, Del Rio et al. DelRio2021 compiled a dataset using English transcripts from 44 earnings calls across nine sectors, recorded in 2020. They meticulously annotated 24 types of entities, including PERSON, ORG, LOC, PRODUCT, EVENT, WORK_OF_ART, LAW, LANGUAGE, DATE, TIME, PERCENT, MONEY, QUANTITY, ORDINAL, CARDINAL, ABBREVIATION, CONTRACTION, ALPHANUMERIC, YEAR, and WEBSITE. The annotation proceeded in three steps: 1) Their custom tool was used to annotate tokens requiring normalization, such as abbreviations, cardinals, ordinals, and contractions; 2) SpaCy 2.3.5333https://spacy.io/ was employed to annotate, for example, ORG and PERSON; and 3) Manual review of annotations and updates were conducted.
Subsequently, in 2022, Loukas et al. Loukas2022 introduced the FINER-139 dataset. This extensive dataset contains 1.1 million sentences tagged with gold XBRL (eXtensible Business Reporting Language) labels. Extracted from reports of publicly traded companies filed with the US Securities and Exchange Commission (SEC) over five years (2016 to 2020), it encompasses the 139 most prevalent XBRL entity types among nearly 6,000 entity types. The authors conducted experiments employing Bi-LSTMs and BERT to evaluate the dataset’s utility and applicability.
Shah et al. Shah2023 , in 2023, introduced a weak supervision NER pipeline for financial texts, named FiNER, and created the FiNER-ORD dataset. This dataset comprises manually annotated financial news articles in English. From a pool of nearly 48,000 news articles, the authors selected 201 and annotated them using the Doccano [^4] https://doccano.github.io/doccano/ software. The annotations focused on generic entities such as PERSON, LOCATION, and ORGANIZATION. Subsequently, the annotated dataset was utilized to validate the efficacy of the developed pipeline.
Similarly, in 2023, Zhang et al. Zhang2023 conducted experiments involving large language models and the recognition of named entities in the Chinese financial domain. They constructed a dataset primarily consisting of text sourced from financial websites, company earnings releases, financial short messages on social media, and financial article summaries, totaling 7,521 sentences. Manual annotation was performed on these sentences, covering six types of generic entities: PERSON_NAME, PRODUCT_NAME, COMPANY_NAME, LOCATION, ORG_NAME, and TIME.
In line with previous research efforts, we have curated a dataset tailored for the NER task. However, our dataset comprises Portuguese text sourced from 384 earnings call transcriptions of 10 banks listed on the Brazilian stock exchange. Employing weak supervision, we annotated the sentences, incorporating both generic and domain-specific labels.
Regarding the fine-tuning of Large Language Models for the NER task, BERTimbau Souza2020 , a Portuguese pre-trained BERT-based model, achieved state-of-the-art performance on the Portuguese MiniHAREM dataset. Similarly, PTT5 Carmo2020 , introduced in 2020, is a Portuguese pre-trained T5 model. After pre-training, it was fine-tuned for various tasks, including NER. In their NER experiments, the authors prefixed the instruction “Recognize Entities” to sentences, as in “Recognize Entities: John lives in New York”. They expected an output following the pattern: “John [Person] lives in [Other] New York [Location]”, where the words in brackets represent the entity types. Experiments were conducted using the Portuguese datasets HAREM and MiniHAREM, widely utilized for NER tasks, resulting in an F1-score of 82%.
In 2021, Finardi et al. Finardi2021 introduced BERTaú, a BERT-based model pre-trained from scratch using data from the Brazilian bank Itaú virtual assistant. This model was fine-tuned for various tasks including Frequent Answer Question Retrieval, Sentiment Analysis, and NER. Their NER dataset, comprising 18,370 manually annotated examples across 16 classes including specific banking entities, yielded an F1-score of 0.877 after 5 epochs of fine-tuning. Additionally, they reported F1-scores of 0.840 for mBERT uncased and 0.853 for BERTimbau base. During inference, a V100 GPU was employed, with an average processing time of 0.01944 seconds per sample.
In 2022, Wang et al. Wang2022instructionner introduced the InstructionNER framework, which leverages the T5-large model to address the Named Entity Recognition (NER) task by converting it from token classification to a sequence-to-sequence (seq2seq) format. Fundamentally, the approach involves feeding the model with a sentence, an instruction, and options. The sentence represents the text in which named entities are to be identified, the instruction guides the model on what action to take, and the options denote the types of entities. The model’s output adheres to the template “entity is a/an type”. The authors employed few-shot NER for fine-tuning, utilizing five datasets distinct from the financial domain. Across their experiments, they achieved a maximum F1 value of 0.954.
Similar to Carmo et al. Carmo2020 , Finardi et al. Finardi2021 , and Wang et al. Wang2022instructionner , our study involved fine-tuning BERT- and T5-based models for the Named Entity Recognition (NER) task. However, our approach differed from Carmo et al. since they introduced PTT5, a pre-trained T5 model from scratch and fine-tuned for different tasks, and we fine-tuned their model and the mT5 model for a domain-specific NER. Additionally, while Finardi et al. pre-trained BERT from scratch and fine-tuned the model for various tasks, we specifically fine-tuned and compared two BERT-based models pre-trained on Portuguese (BERTimbau) and multilingual (mBERT) datasets exclusively for NER. Moreover, our focus was on domain-specific entities such as profit, revenue, net income, and cash flow, whereas their focus was on entities related to customer services.
In contrast to the approach of Wang et al. Wang2022instructionner , who employed the large version of T5, we fine-tuned two T5-based models with fewer parameters pre-trained on Portuguese (PTT5 base version) and multilingual (mT5 small version) datasets. Similarly, comparable to the methodologies of Carmo et al. Carmo2020 and Wang et al. Wang2022instructionner , we reformulated the NER task from token classification to seq2seq. However, our approach diverges from theirs in that we fine-tuned the models using a sentence as input and generated the same sentence as output, with entities and their labels annotated within the sentence.
@@ 4 Methodology
To evaluate Transformer-based models for the NER task, we constructed a dataset from earnings calls of Brazilian banks by extracting and annotating the sentences. Subsequently, we fine-tuned four models and analyzed their results. Below, we outline the main steps to conduct the experiments. The scripts and datasets are available in a public repository [^5] https://github.com/rsabilio/NerEval-BrazilianCorporateTranscriptions.
4.1 Step 1: Download PDF files of Earnings Call
In Step 1, we verified the Comissão de Valores Mobiliários (CVM) open data CVM2023 to identify active banks. We found 29 banks and, subsequently, we visited their Investor Relations (IR) website to check if they provided their earnings call transcriptions. In positive cases, we compiled a list containing the bank’s data such as: name, ticker (code of negotiation on the Brazilian Stock Exchange), the year when the bank started to provide the transcriptions, and the URL of the IR website. Using this list, we developed a crawler to download the PDF files of each transcription published until the second quarter of 2023.
Table 1 presents the ten selected banks along with their name, ticker, the year when the bank started providing the transcriptions, and the majority ownership indicating whether the bank is controlled by the government or by private individuals or entities. We observe that two banks are controlled by the government: Banco do Brasil (Federal Government) and Banco Estado do Rio Grande do Sul (State of Rio Grande do Sul), while eight are private banks. Notably, Banco do Brasil was the first bank to provide its transcriptions in 2006. It is important to mention that, as of February 2024, the publication of earnings call transcriptions is not mandatory. Hence, while other banks may provide video or audio of the earnings calls, some of the selected banks may not provide transcriptions for specific quarters.
@@ 4.1 Step 1: Download PDF files of Earnings Call
4.2 Step 2: Extract and Annotate texts from PDF
In this step, we built a list of labels based on our objective which was to recognize entities from the financial domain present in the bank’s earnings call such as: Percentage, Money, Period (quarter, semester), Profit, Revenue, and bank’s products. We also built a gazetteer with labels and examples of entities. This gazetteer was used to develop a script based on Skweak v0.3.3 package Lison2021 to recognize those entities and annotate them.
To extract the texts from the PDF files, we used the PDFPlumber [^6] https://github.com/jsvine/pdfplumber v0.10.3 package. In the extraction process, we used NLTK [^7] https://www.nltk.org/ v3.8 for sentence tokenization and the sentences were saved in a CSV file. The sentences in this file were annotated by Skweak.
Skweak provides mechanisms based on regular expressions, functions (heuristics), and gazetteers Lison2021 . We implemented regular expressions to annotate year, quarter, and semester and functions to annotate percentage and money. We used a JSON file containing the gazetteer. Skweak annotated each sentence using the regular expressions, the functions, and the gazetteer. At the end, it used a Hidden Markov Model to aggregate different annotations for the same entity resulting in a JSON structure containing a collection of sentences and their named entities (entity, label, start and end positions).
4.3 Step 3: Select and fine-tune models
We searched on the Hugging Face platform for models provided by research groups or companies, specifically focusing on those pre-trained either exclusively in Portuguese or on multilingual datasets. We identified BERTimbau and PTT5 as models pre-trained in Portuguese, and mBERT and mT5 as multilingual models. While reviewing the available versions on Hugging Face, we opted for versions recommended by the original authors and with a comparable number of parameters to facilitate result comparison. Additionally, we included the smallest versions due to infrastructure constraints, despite the expectation of better results with larger models Devlin2019 ; Raffel2020 .
Table 2: Selected Models
Architecture Model Language Version #Parameters
BERT BERTimbau Portuguese Base 110M
BERT mBERT Multilingual Base 110M
T5 PTT5 Portuguese Base 220M
T5 mT5 Multilingual Small 300M
Before the fine-tuning, we separated the annotated sentences into three files: train.csv, val.csv, and test.csv. We follow the proportion: 70%, 20%, and 10%. Besides, we studied the models’ input format and their tokenizers. After that, we developed scripts using Google Colaboratory (Colab) platform, Python v3.10, and Hugging Face libraries such as Transformers v4.35.
We prepared the data developing different scripts for each architecture due to its characteristics. For example, while BERT-based models expect a sentence and a sequence of labels as inputs, T5-based models expect a sentence and an expected text (text that would be generated). For both architectures, we used the Beginning, Inside, and Outside (BIO) tagging scheme Ramshaw1995-BIO .
4.4 Step 4: Evaluation of the models results
We saved the results of each model in separate CSV files containing: i) sentence ID, a sequential number to uniquely identify each sequence; ii) annotated token, a token obtained from Skweak annotation; iii) annotated label, the label attributed by Skweak to a token; iv) model token, the token identified by the model; and v) model label, the label attributed by the model. We processed each file assigning a category to each register according to the Fifth Message Understanding Conference (MUC-5) Chinchor1993-muc :
  • 1.
  • 2.
  • 3.
  • 4.
We calculated Primary and Secondary metrics based on the categories, where lower values indicate better performance. The Primary metric “Error per response fill”, adapted from MUC-5 Chinchor1993-muc , is expressed by Eq. 1.
$Error\ per\ response\ fill=\frac{(INC+MIS+SPU)}{(COR+INC+MIS+SPU)}$ (1)
$Undergeneration=\frac{MIS}{COR+INC+MIS}$ (2)
$Overgeneration=\frac{SPU}{COR+INC+SPU}$ (3)
$Substitution=\frac{INC}{COR+INC}$ (4)
5 Dataset development
This section presents the dataset development, which comprises the acquisition of the transcription files and the annotation of the extracted sentences.
5.1 Transcriptions acquisition
@@
We acquired 384 files published between 2006 and 2023, spanning 18 years, from the bank’s Investor Relation website (Table 1). The distribution of transcriptions per year and bank is visualized in Fig. 1. For instance, in 2022, we obtained 34 transcriptions from 9 banks, whereas only one was available from one bank in 2006. Such variations arise due to the diverse starting points for publication among the banks, occurring in different epochs, whether in a specific year or quarter.
Figure 1: Number of transcriptions per year and bank
@@
After the transcriptions download, we extracted their texts breaking them into sentences. At the end of this process, we obtained 118,411 sentences. We analyzed those sentences observing their length regarding the number of words without punctuation and sentence duplicates.
@@
We observed sentences ranging from 1 to 262 words in length. Sentences with up to 4 words were filtered for further analysis, resulting in 9,103 sentences. Upon analysis, it was determined that these short sentences, primarily greetings or phrases such as “Por favor, aguardem.” (“Please, wait.”), “Obrigado.” (“Thank you.”), and “Muito obrigado a todos.” (“Thank you all”), would not significantly contribute to the domain. Additionally, 10,833 duplicated sentences were identified, including shorter and longer sentences such as “Ótimo.” (“Great.”) and “A sessão de perguntas e respostas está encerrada.” (“The question and answer session is closed.”). To enhance the quality of the dataset for fine-tuning models, we chose to exclude both duplicated sentences and those with fewer than 4 words.
@@
After this step, we amassed a corpus comprising 103,340 sentences, totaling 2,575,053 words. To elucidate the distribution of sentence lengths across different temporal segments, we conducted an analysis employing a histogram, as depicted in Fig. 2. This graphical representation illustrates the relationship between sentence length, denoted on the x-axis, and the corresponding frequency of sentences, depicted on the y-axis, for each calendar year. For instance, our scrutiny of transcriptions from the year 2022, totaling 34 in the count, revealed a spectrum of sentence lengths ranging from 5 to 236 words. Notably, within this range, a significant portion of 9,146 sentences exhibited lengths between 5 and 28 words.
@@
5.2 Data annotation
In the annotation process using weak supervision, we applied 23 labels, encompassing both generic labels, such as PERCENTUAL (Percentage), COMPANY, MONEY, YEAR, and ORG, as well as domain-specific labels including CARTEIRA (Portfolio), CLIENTE (Client), RISCO (Risk), PROVISAO (Reserve), DESPESA (Expense), PRODUTO (Product), QUARTER, CONDICOES_MACROECONOMICAS (Macroeconomic Conditions), INDICADOR_LIQUIDEZ (Liquidity Indicator), INDICADOR_EFICIENCIA (Efficiency Indicator), RECEITA (Revenue), BALANCO_PATRIMONIAL (Balance Sheet), SEMESTER, LUCRO (Profit), INDICADOR_RENTABILIDADE (Profitability Indicator), PROVENTO (Dividends), INDICADOR_VALUATION (Valuation Indicator), and RESULTADOS (Results) (Table 3). The incorporation of these domain-specific labels distinguishes our approach from related work that typically relies on generic labels, even when dealing with datasets based on financial texts.
After running Skweak, a total of 119,517 annotations were generated from 57,933 sentences, resulting in an average of 2 annotations per sentence. From the total of sentences, the 45,407 sentences without annotations were filtered and stored in a separate CSV file. The distribution of entities per label is detailed in Table 3, arranged in descending order. Notably, the label “PERCENTUAL” (Percentage) exhibited the highest frequency with 26,934 occurrences, whereas “RESULTADO” (Result) was the least frequent, with only 81 instances.

### v2

@@ 3.2 T5-based models: PTT5 and mT5
The Text-to-Text Transfer Transformer (T5) model is based on the Transformer encoder-decoder architecture and uses SentencePiece as its tokenizer [18]. To use T5, we need to map our problem to a text-to-text format, i.e. taking text as input and generating text as output [18]. T5 was pre-trained using “Colossal Clean Crawled Corpus” (C4) and it was evaluated on a wide variety of NLP tasks such as question answering, document summarization, and sentiment classification [18]. Instead of using a fixed embedding for each position like BERT, T5 uses relative position embeddings.
@@
Similar to BERT, T5 has also served as a base for other models. Carmo et al. [23] introduced, in 2020, a model based on T5 architecture and pre-trained from scratch using BrWaC (Brazilian Web as Corpus) dataset. The authors named this model PTT5 and released three versions on Hugging Face: Small (60M parameters), Base (220M parameters), and Large (740M parameters). They also recommend the use of the Base version.
@@
A multilingual version of T5 was introduced by Xue et al. [24] in 2021. The model, named mT5, was pre-trained using a multilingual version of C4 (mC4 - 101 languages). mT5 was released in five versions: mT5-Small (300M parameters), mT5-Base (580M parameters), mT5-Large (1.2B parameters), mT5-XL (3.7B parameters), and mT5-XXL (13B parameters).
@@
T5 models can process longer inputs than BERT models. However, longer inputs may still lead to increased computational cost and memory usage during training and inference. We confirmed that PTT5 and mT5 provided on Hugging Face have an input length of 512.
@@ 4 Methodology
We conducted this work in three main steps, as shown in Fig. 1. In the first step (Dataset), we constructed a dataset from earnings calls of Brazilian banks by extracting and annotating the sentences. Subsequently, in the Fine-tuning step, we fine-tuned four models. Finally, we analyzed their inference results in the Evaluation step. These steps are outlined in the following subsections. The scripts and datasets are available in a public repository [^5] https://github.com/rsabilio/NerEval-BrazilianCorporateTranscripts.
Figure 1: Overview of the main methodological steps
4.1 Dataset development
Earnings calls are conferences held by publicly traded companies to present their financial performance and results to stakeholders, including investors, analysts, and the media. These calls typically feature presentations by the company’s executives, followed by a Question & Answer session where analysts and investors can ask questions. The discussions often cover key financial statements such as the balance sheet, income statement, and cash flow statement. Some companies choose to transcribe these calls and make the transcripts publicly available. These transcripts are particularly interesting for study because they capture the nuances and informalities of spoken language, which can differ significantly from written text. This variability presents challenges for automatic information extraction. For example, the phrase ‘O lucro foi de 900 milhões de reais” (The profit was BRL 900 million) might be transcribed as “O lucro foi R\$900 milhões”, “O lucro foi de R\$ 900 MM”, or simply “O lucro foi de 900 milhões de reais”. Given these challenges, achieving high performance in Named Entity Recognition (NER) tasks by language models is crucial.
The following subsections present the development of the Brazilian Financial NER Dataset (BraFiNER), which comprises the acquisition of the transcript files and the annotation of the extracted sentences.
4.1.1 Download of PDF files of Earnings Call
We reviewed the Comissão de Valores Mobiliários (CVM) [^6] The Comissão de Valores Mobiliários (CVM), akin to the U.S. Securities and Exchange Commission (SEC), is responsible for the development, regulation, and supervision of the Brazilian Securities Market. The CVM’s mission is to facilitate capital raising for companies, safeguard investor interests, and ensure comprehensive disclosure of information regarding issuers and their securities. open data [31] and identified 29 active banks. We then visited their Investor Relations (IR) websites to check for available earnings call transcripts. For banks providing the PDF files, we compiled a list including the bank’s name, ticker (code of negotiation on the Brazilian Stock Exchange), start year for transcripts, and IR website URL. Using this list, we developed a crawler to download the PDF files published up to the second quarter of 2023.
Table 1 presents the ten selected banks along with their name, ticker, the year when the bank started providing the transcripts, and the majority ownership indicating whether the bank is controlled by the government or by private individuals or entities. We observe that two banks are controlled by the government: Banco do Brasil (Federal Government) and Banco do Estado do Rio Grande do Sul (State of Rio Grande do Sul), while eight are private banks. Notably, Banco do Brasil was the first bank to provide its transcripts in 2006. It is important to mention that, as of June 2024, the publication of earnings call transcripts is not mandatory. Hence, while other banks may provide video or audio of the earnings calls, some of the selected banks may not provide transcripts for specific quarters.
@@ 4.1.1 Download of PDF files of Earnings Call
With the crawler, we acquired 384 files published between 2006 and 2023, spanning 18 years. The distribution of transcripts per year and bank is visualized in Fig. 2. For instance, in 2022, we obtained 34 transcripts from 9 banks, whereas only one was available from one bank in 2006. Such variations arise due to the diverse starting points for publication among the banks, occurring in different epochs, whether in a specific year or quarter.
Figure 2: Number of transcripts per year and bank
@@
After the transcripts download, we extracted the sentences from the PDF files using the PDFPlumber [^7] https://github.com/jsvine/pdfplumber v0.10.3 package. Subsequently, we employed NLTK [^8] https://www.nltk.org/ v3.8 for sentence tokenization. Finally, all segmented sentences were compiled and stored in a CSV file for further processing and annotation. At the end of this process, we obtained 118,411 sentences. We analyzed those sentences observing their length regarding the number of words without punctuation and sentence duplicates.
@@
We observed sentences ranging from 1 to 262 words in length. We filtered 9,103 sentences with up to 4 words, and upon analysis, determined that these short sentences, primarily greetings or phrases such as “Por favor, aguardem.” (“Please, wait.”), “Obrigado.” (“Thank you.”), and “Muito obrigado a todos.” (“Thank you all.”), would not significantly contribute to the domain. Additionally, we identified 10,833 duplicated sentences, including shorter and longer sentences such as “Ótimo.” (“Great.”) and “A sessão de perguntas e respostas está encerrada.” (“The question and answer session is closed.”). To enhance the quality of the dataset for fine-tuning models, we excluded both duplicated sentences and those with fewer than 4 words.
@@
After this step, we amassed a corpus comprising 103,340 sentences, totaling 2,575,053 words. To elucidate the distribution of sentence lengths across different temporal segments, we plotted histograms (Fig. 3) with sentence length on the x-axis and the corresponding frequency of sentences on the y-axis, for each calendar year. For instance, our analysis of transcripts from 2022, totaling 34, revealed a spectrum of sentence lengths ranging from 5 to 236 words. Notably, within this range, a significant portion of 9,146 sentences exhibited lengths between 5 and 28 words.
@@
4.1.2 Data annotation
We used a weakly supervised approach based on regular expressions, heuristics, and gazetteers [32] to annotate the entities present in the extracted sentences. As an initial step, we defined a list of 23 labels, encompassing both generic labels, such as PERCENTUAL (Percentage), COMPANY, MONEY, YEAR, and ORG, as well as domain-specific labels including CARTEIRA (Portfolio), CLIENTE (Client), RISCO (Risk), PROVISAO (Reserve), DESPESA (Expense), PRODUTO (Product), CONDICOES_MACROECONOMICAS (Macroeconomic Conditions), INDICADOR_LIQUIDEZ (Liquidity Indicator), INDICADOR_EFICIENCIA (Efficiency Indicator), QUARTER, LUCRO (Profit), BALANCO_PATRIMONIAL (Balance Sheet), SEMESTER, INDICADOR_RENTABILIDADE (Profitability Indicator), RECEITA (Revenue), PROVENTO (Dividends), INDICADOR_VALUATION (Valuation Indicator), and RESULTADOS (Results). The incorporation of these domain-specific labels distinguishes our approach from related work that typically relies on generic labels, even when dealing with datasets based on financial texts.
From that list, we developed a script using the Skweak package v0.3.3 [32] to recognize those entities and annotate them. Skweak provides mechanisms based on regular expressions, functions (heuristics), and gazetteers. We implemented regular expressions to annotate year, quarter, and semester, functions to annotate percentage and money, and gazetteers to annotate the other entities. At the end of the annotation process, Skweak used a Hidden Markov Model to aggregate different annotations for the same entity resulting in a JSON structure containing a collection of sentences and their named entities (entity, label, start and end positions).
After running Skweak, a total of 119,517 annotations were generated from 57,933 sentences, resulting in an average of 2 annotations per sentence. From the total of sentences, the 45,407 sentences without annotations were filtered and stored in a separate CSV file. The distribution of entities per label is detailed in Table 2, arranged in descending order. Notably, the label “PERCENTUAL” (Percentage) exhibited the highest frequency with 26,934 occurrences, whereas “RESULTADO” (Result) was the least frequent, with 81 instances.